OpenStudy (anonymous):

evaluate radical -36m^2

OpenStudy (anonymous):

\[\sqrt{-36m^2}=\sqrt{36}\sqrt{m^2}\sqrt{-1}=6mi\]

OpenStudy (anonymous):

i assume the minus sign was under the radical

OpenStudy (anonymous):

yes, thank you

OpenStudy (anonymous):

yw